OpenStudy (anonymous):

HELPPPPPPP!!! How do you solve this?!?!?! 7.2 * 10^8 dm/min^2 = ? m/s^2 I GOT: 2.592 * 10^10 IS THAT RIGHT! HELP :((((

OpenStudy (anonymous):

(7.2 * 10^8)/ (3600 * 10) will give u an answer in m/s^2

OpenStudy (xishem):

Dimensional analysis: \[7.2 \times 10^8 \frac{dm}{\min^2} \cdot \frac{1m}{10dm} \cdot \frac{1\min^2}{3600s^2}\]